How long will it take for an investment to double in value, if the annual interest rate is 10% and interest is compounded every 6 months?
Answer by lwsshak3(11628)

How long will it take for an investment to double in value, if the annual interest rate is 10% and interest is compounded every 6 months
**
compound interest formula: A=P(1+i)^n, A=value of investment after n periods, P=initial investment, i=interest rate per period, n=number of periods.
For given problem:A/P=(1+i)^n,A/P=2, i=.10/2=.05
2=(1+.05)^n
2=(1.05)^n
take log of both sides
log2=nlog(1.05)
n=log2/log(1.05)
n=14.2 6-mo periods or 7.1 years